# Cocoa: Co-Planning and Co-Execution with AI Agents

### v3

@@ Abstract.
Human collaboration benefits from continuous coordination— planning, delegating tasks, sharing progress, and adjusting objectives—to align on shared goals. However, agentic AI systems often limit users to previewing or reviewing an agent’s plans for fully autonomous execution. While this may be useful for confirmation and correction, it does not support deeper collaboration between humans and AI agents. We present Cocoa, a system that introduces a novel design pattern—interactive plans—for collaborating with an AI agent on complex, multi-step tasks. Informed by a formative study ($n=9$), Cocoa builds on interaction designs from computational notebooks and document editors to support flexible delegation of agency through Co-planning and Co-execution, where users collaboratively compose and execute plans with an Agent. Using scientific research as a sample domain, our lab ($n=16$) and field deployment ($n=7$) studies found that Cocoa improved agent steerability without sacrificing ease-of-use compared to a strong chat baseline. Additionally, researchers valued Cocoa for real-world projects and saw the interleaving of co-planning and co-execution as an effective novel paradigm for human-AI collaboration.
copyright: acmcopyrightjournalyear: 2025doi: xxxxxxx.xxxxxxxconference: arXiv; April; 2025price: 15.00isbn: xx-x-xxx-xxxx-xxxxx
@@ 1. Introduction
Figure 2. An overview of the Cocoa user interface. An interactive plan (A) affords human-AI co-planning and co-execution: a researcher and the AI agent can collaboratively edit the plan in the document and execute the plan steps, similar to executing code cells in a computational notebook. Steps can be assigned to the AI agent (B) or the researcher (C). The researcher can freely edit the AI agent’s outputs in an interactive sidebar (D), such as adding relevant papers that the agent did not find (E) to help steer the agent with their feedback and expertise. In this example, the first three steps of a plan to summarize methods for human feedback elicitation have already been executed, and the agent is requesting guidance from the user in the next step.
[…]
@@
An overview of the Cocoa user interface, which resembles a document editor. An interactive plan affords human-AI co-planning and co-execution: a researcher and the AI agent can collaboratively edit the plan in the document and execute the plan steps, similar to executing code cells in a computational notebook. Steps can be assigned to the AI agent or the researcher. The researcher can freely edit the AI agent’s outputs in an interactive sidebar, such as adding relevant papers that the agent did not find, to help steer the agent with their feedback and expertise. In the example shown, the first three steps of a plan to summarize methods for human feedback elicitation have already been executed, and the agent is requesting guidance from the user in the next step.
@@
Since the advent of personal computing, researchers and practitioners in artificial intelligence (AI) and human-computer interaction (HCI) have set sights on developing intelligent AI agents that can help perform everyday computer-based tasks on our behalf (Maes, 1994; Maes and Wexelblat, 1996; Lieberman, 1997; Horvitz, 1999). Recent advancements in large language models (LLMs) and agent frameworks involving reasoning, planning, memory, and tool use (Yao et al., 2022b; Sumers et al., 2023; Schick et al., 2023; Wang et al., 2023; Kim et al., 2024) have accelerated the development of fully autonomous AI agents that can conduct complex tasks with varying degrees of success. These tasks include shopping online (Yao et al., 2022a), writing software (Yang et al., 2024; Jimenez et al., 2023; Wu, 2024), performing “deep research” to assemble a report based on web sources (Google, 2025; OpenAI, 2025a; Perplexity, 2025), and general computer-based tasks (Anthropic, 2024; OpenAI, 2025b). By contrast, turn-based and direct manipulation LLM systems that shift the burden of task planning and progress monitoring to users—such as chat or node-based interfaces (Suh et al., 2023; Liu et al., 2024a; Wu et al., 2022; Pu et al., 2024)—have become popular in HCI in recent years for offering users greater agency and control.
While some systems have attempted to balance AI and human agency by allowing users to review an agent’s reasoning chains after execution for verification (OpenAI, 2024; Guo et al., 2025), or to edit and confirm its plans before execution (Google, 2025; OpenAI, 2025b), few works have explored how to deeply involve and collaborate with users during iterative planning and execution. We argue that effective human interaction with agents remains essential for agents to operate effectively and safely in the real world (Kapoor et al., 2024; Shi et al., 2024), such that human-AI teams can accomplish tasks that cannot easily be completed by humans or AI alone (Bansal et al., 2021). Realizing this vision demand deeper explorations into strategies for fostering synergistic collaboration between human users and AI agents (Kapoor et al., 2024; Collins et al., 2024; Ibrahim et al., 2024).
More concretely, this lack of human-AI collaboration in task planning and execution can lead to several issues. First, we cannot steer the agent with our expertise and worldly understanding. In domains where agents can not yet produce a quality plan and the ability to steer agents towards one, agents can easily veer off track, wasting resources without achieving meaningful results (Valmeekam et al., 2023; Husain et al., 2025). When executing the plan, human guidance can significantly boost agent performance—Shi et al. (Shi et al., 2024) found that even simple feedback from programmers to an agent for solving Olympiad-style programming problems increased agent accuracy from 0% to over 85%. Second, and perhaps more importantly, removing human agency in favor of AI agency can pose heightened safety risks, disempower us to think critically and be creative, and harm our well-being more generally (Chan et al., 2023; Bennett et al., 2023; Kulveit et al., 2025; Collins et al., 2024). More fundamentally, for many nuanced, subjective, and high-stakes tasks, human input must be considered for AI agents to be successful and aligned with users’ personal needs and goals (Cheong et al., 2024; Feng et al., 2024; Terry et al., 2023).
@@
In this work, we introduce Cocoa, a novel system for Co-planning and Co-execution with AI Agents. We chose scientific research tasks as a sample domain due to the complex, multi-step, and information-dense workflows researchers often engage in throughout their work (Kang et al., 2022b; Fok et al., 2022, 2023; Kang et al., 2022c; Lee et al., 2024b). Cocoa embeds AI agents into a document editor—a common environment many researchers use to keep track of tasks and progress for their research projects. Taking inspiration from computational notebook interfaces where data scientists compose program instructions and manage their execution and outputs, we introduce a new interaction design pattern for AI agents called interactive plans. Interactive plans orchestrate actions between a human user and an AI agent and enable flexible delegation of human and AI agency. As a result, Cocoa supports human-AI co-planning: the agent, when invoked in the document, will first propose a plan of action that is interactive and seamlessly integrates into the document. The user can then edit the plan steps through familiar interactions that mirror typical document editing and assign steps to the agent or themselves. Cocoa also supports human-AI co-execution: drawing inspiration from the design of computational notebooks, the interactive plan allows the user and the agent to collaboratively complete one step at a time or re-execute steps as desired. The user can interactively refine the agent’s intermediate outputs and also manually take over steps themselves to steer the workflow in a more effective direction. Most importantly, Cocoa interleaves co-planning and co-execution, such that users can smoothly transition between the two and modify their plans based on outputs from execution, and vice versa.
@@
    A formative study with 9 researchers with a focus around their real-world project documents that uncovered needs and opportunities to better support planning and execution.
@@
    Cocoa, an interactive system that implements a new design pattern—interactive plans—in a document editor for researchers to engage in co-planning and co-execution with an AI agent.
@@
    A user evaluation of Cocoa with 16 researchers, where we found that interactive plans with novel interaction affordances enabled users to better steer the AI agent without sacrificing ease of use when compared to a conversational chat baseline.
@@
    A 7-day deployment study of Cocoa with 7 researchers that offers insights into how interactive plans can provide value to researchers in their day-to-day work.
@@ 2.1. Planning and Interactivity in LLM Agents
Prior work has identified core components in LLM agent architectures to consist of memory, reasoning, planning, and tool use (Yao et al., 2022b; Zhou et al., 2023; Liu et al., 2024b; Sumers et al., 2023; Wang et al., 2023; Kim et al., 2024). Central to LLM agents’ longitudinal operations is long-term planning, which demands capable reasoning abilities (Hao et al., 2023) and thus is often implemented by chain-of-thought (CoT) [^1] We use CoT as an umbrella term encompassing chain-of-thought, trees-of-thought (Yao et al., 2023; Zhuang et al., 2023), and related methods. (Wei et al., 2022; Qin et al., 2023; Yao et al., 2022b; OpenAI, 2024). CoT provides a series of intermediate reasoning steps as exemplars in prompting to boost performance on complex reasoning tasks (Wei et al., 2022; OpenAI, 2024). CoT is crucial for facilitating task decomposition and therefore LLM agents’ planning capabilities (Qin et al., 2023; Chen et al., 2023). However, even with CoT, critical investigations of LLMs’ abilities to autonomously generate executable plans reveal limited success across diverse domains (Valmeekam et al., 2023). Indeed, although CoT can help improve model planning for tasks with well-defined, objective solutions that provide clear signals for reinforcement learning—e.g., numerical, tabular, and knowledge-based reasoning (Kim et al., 2024)—it may not be so effective in domains that require expert tacit knowledge to navigate ambiguous problem spaces with no single right answer (Hao et al., 2023). Unlike model properties that show empirical improvement through scaling laws, limitations of planning in these domains may not resolve with scale alone as 1) tacit knowledge is not well-documented in training data and is thus difficult for a model to robustly learn (Cheong et al., 2024; Feng et al., 2023; Zhang et al., 2024), and 2) there may be no “correct” workflow for CoT to follow and verify the correctness post-hoc (Jirotka et al., 2013). Scientific research is one such domain (Zhang et al., 2024).
@@
In light of this, recent work has recognized the need to interactively incorporate user feedback for improving LLM agents’ planning capabilities and beyond (Valmeekam et al., 2023; Yao et al., 2022b; Kazemitabaar et al., 2024; Lawley and MacLellan, 2024; Shao et al., 2024; OpenAI, 2025b), particularly within scientific discovery (Majumder et al., 2024; Gottweis et al., 2025). For example, Lawley and MacLellan (Lawley and MacLellan, 2024) architected an approach where user interaction is used to guide the model in planning for unseen tasks on-the-fly using a hierarchical network of smaller actions. OpenAI’s Operator (OpenAI, 2025b) allows users to “take control” of the agent to perform tasks manually. Yet, interactive techniques in this area are still nascent, despite the new challenges identified for human-agent communication (Bansal et al., 2024). Interfaces for LLM assistants (e.g., AgentGPT (Reworkd, 2024), Devin (Wu, 2024)) have been primarily limited to chat interfaces targeted at monitoring agent activity rather than empowering the user to proactively collaborate with the agent. An emerging body of work has started to experiment with more interactive techniques for agents. Kazemitabaar et al. (Kazemitabaar et al., 2024) developed interfaces for data analysts to edit an execution plan of an LLM to provide more control points for steering behavior, while Google’s Deep Research (Google, 2025) can re-generate a plan if the user provides feedback via prompting. However, these interactions are corrective rather than collaborative—that is, users would typically engage in these interactions to correct agent behavior post-hoc rather than proactively iterate back-and-forth with the agent at multiple points in the workflow. We target the latter in our work to elevate human feedback to beyond simply a corrective mechanism. That is, our goal is to enable interactive agents that effectively elicit and incorporate key guidance from the user through co-created human-AI representations.
@@ 2.2. Computational Notebooks
The computational notebook is an interactive paradigm that organizes program imperatives, input data, and rich outputs (e.g., tables, visualizations, interactive widgets) into linearly arranged cells (Chattopadhyay et al., 2020; Lau et al., 2020), reflecting Donald Knuth’s early visions of literate programming (Knuth, 1984). Computational notebooks are a well-studied paradigm in HCI (e.g., (Choi et al., 2023; Ayobi et al., 2023; Zheng et al., 2022; Lin et al., 2023; Chattopadhyay et al., 2020; O’Leary et al., 2023)), with most prior work focusing on how data scientists leverage them in their workflows.
In our work, we take inspiration from design decisions made for computational notebooks because of the many parallels between workflows in notebooks and mixed-initiative task completion systems (Horvitz, 1999; Grunde-McLaughlin et al., 2023; Wu et al., 2023). For instance, data scientists often break down a high-level data analysis task into executable cells in computational notebooks (Kery et al., 2018; Rule et al., 2018). As they execute each cell and inspect outputs, they can verify the quality and sensibility of outputs before moving forward in their analysis, or go back to modify existing code cells before execution to interleave between programming, execution, and output examination (Kery et al., 2020).
By drawing these parallels, we can reveal new opportunities to design interactive workflows and interfaces for LLM agents. Just like how a data scientist may easily add new cells in a notebook or reconfigure existing cells to adapt their analysis plan, a user may interactively edit an agent’s plan of execution to better steer the agent in productive directions. Furthermore, an agentic system may simultaneously be more usable and resource-efficient if a user could direct an agent to iterate on a subtask to improve its output before continuing onto subtasks dependent on that output. This also demands new ways of viewing and editing agent outputs on subtasks, which increases opportunities for human input over existing approaches such as simply logging agent actions (AutoGPT, 2024; Reworkd, 2024; OpenAI Platform, 2024). Our work exploits these parallels to contribute new design patterns for steering agentic systems.

### v4

@@ Abstract.
As AI agents take on increasingly long-running tasks involving sophisticated planning and execution, there is a corresponding need for novel interaction designs that enable deeper human-agent collaboration. However, most prior works leverage human interaction to fix “autonomous” workflows that have yet to become fully autonomous or rigidly treat planning and execution as separate stages. Based on a formative study with 9 researchers using AI to support their work, we propose a design that affords greater flexibility in collaboration, so that users can 1) delegate agency to the user or agent via a collaborative plan where individual steps can be assigned; and 2) interleave planning and execution so that plans can adjust after partial execution. We introduce Cocoa, a system that takes design inspiration from computational notebooks to support complex research tasks. A lab study ($n=16$) found that Cocoa enabled steerability without sacrificing ease-of-use, and a week-long field deployment ($n=7$) showed how researchers collaborated with Cocoa to accomplish real-world tasks.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791673isbn: 979-8-4007-2278-3/2026/04
@@ 1. Introduction
Figure 2. An overview of the Cocoa user interface. An interactive plan (A) affords human-agent co-planning and co-execution: a researcher and the AI agent can collaboratively edit the plan in the document and execute the plan steps, similar to executing code cells in a computational notebook. Steps can be assigned to the AI agent (B) or the researcher (C). The researcher can freely edit the AI agent’s outputs in an interactive sidebar (D), such as adding relevant papers that the agent did not find (E) to help steer the agent with their feedback and expertise. In this example, the first three steps of a plan to summarize methods for human feedback elicitation have already been executed, and the agent is requesting guidance from the user in the next step.
@@
An overview of the Cocoa user interface, which resembles a document editor. An interactive plan affords human-agent co-planning and co-execution: a researcher and the AI agent can collaboratively edit the plan in the document and execute the plan steps, similar to executing code cells in a computational notebook. Steps can be assigned to the AI agent or the researcher. The researcher can freely edit the AI agent’s outputs in an interactive sidebar, such as adding relevant papers that the agent did not find, to help steer the agent with their feedback and expertise. In the example shown, the first three steps of a plan to summarize methods for human feedback elicitation have already been executed, and the agent is requesting guidance from the user in the next step.
@@
Since the advent of personal computing, researchers and practitioners in artificial intelligence (AI) and human-computer interaction (HCI) have set sights on developing AI agents that can collaborate with humans to perform complex computer-based tasks (Maes, 1994; Maes and Wexelblat, 1996; Lieberman, 1997; Horvitz, 1999). For example, paradigms such as mixed-initiative interaction envisioned fluid human-AI collaboration with user-led direct data manipulation combined with AI-led autonomous task execution (Horvitz, 1999). Recent advancements in large language models (LLMs) and agent frameworks involving reasoning, planning, memory, and tool use (Yao et al., 2022b; Sumers et al., 2023; Schick et al., 2023; Wang et al., 2023; Kim et al., 2024) have accelerated the development of fully autonomous AI agents to tackle complex tasks, including shopping online (Yao et al., 2022a), writing software (Yang et al., 2024; Jimenez et al., 2023; Wu, 2024), performing “deep research” to assemble a report based on web sources (Google, 2025; OpenAI, 2025b; Perplexity, 2025), and general computer use (Anthropic, 2024; OpenAI, 2025a; Shaikh et al., 2025). Despite the push towards full autonomy, recent work has recognized the importance of deeper human-agent collaboration for not only bolstering agent performance on complex tasks (Shi et al., 2024), but also improving accessibility (Peng et al., 2025), user control and trust (Mozannar et al., 2025; Huq et al., 2025), safety (OpenAI, 2025a), and user agency more generally (Feng et al., 2025), leading to the development of more collaborative human-agent systems, such as for web browsing (Peng et al., 2025; Huq et al., 2025; Mozannar et al., 2025) and general productivity (Shao et al., 2024; OpenAI, 2025a).
However, many of these systems incorporate prescribed collaborative workflows that may limit their utility and steerability. Some works use agent-guided workflows: the agent dictates what input to elicit from the user, as well as when and how the elicitation happens (Peng et al., 2025; Huq et al., 2025; OpenAI, 2025a; Mozannar et al., 2025; Shao et al., 2024). Users may intervene by manually interrupting or taking control of the agent (Mozannar et al., 2025; OpenAI, 2025a, b), but these efforts are reactive and rely on persistent awareness of where the agent has erred. On the other hand, researchers have developed systems with user-guided workflows, where the user is in full control of task planning, while the agent executes scoped subtasks (He et al., 2025; Kazemitabaar et al., 2024; Suh et al., 2023; Angert et al., 2023). Whether agent- or user-guided, these systems encode a particular configuration of human-agent collaboration—there are few affordances to flexibly delegate agency by altering who is guiding based on environmental context or user preferences (Satyanarayan, 2024).
Separately, researchers and industry practitioners have realized the value of decoupling human-agent planning and execution to improve control and avoid fruitless execution (Huyen, 2025; He et al., 2025; Kazemitabaar et al., 2024). However, completely isolating the two into rigid stages can lead to poor results as well—plans often need to be iteratively adapted based on results from completed steps (Mozannar et al., 2025). In contrast, fully autonomous frameworks like ReAct (Yao et al., 2022b) interleaves planning and execution, and has shown benefits in allowing fully autonomous AI agents to readjust their plans after seeing partial execution results. We thus see a new design opportunity for human-agent interaction and ask: what might flexible configurations of agency in human-agent collaboration across planning and execution look like, and how will users make use of them in their own work?
In this work, we explore a novel instantiation of mixed-initiative interaction (Horvitz, 1999) for explicitly and flexibly delegating work in human-agent hybrid workflows by interleaving collaborative planning and execution. We conduct this exploration in the domain of scientific research, a challenging domain for human-agent collaboration due to the prevalence of knowledge-intensive workflows and nuanced decisions made with domain-specific knowledge. Existing agentic research tools primarily focus on supporting human-agent collaboration either before or after task execution. One group of systems elicits user intent and preferences before execution (OpenAI, 2025b; Google, 2025; Perplexity, 2025) (e.g., asking clarification questions about users’ goals (OpenAI, 2025b), allowing plan review and modification (Google, 2025; Mozannar et al., 2025)). Conversely, another group helps users verify and correct the agent after it has generated an output after executing its plan (OpenAI, 2024a; Guo et al., 2025). In practice, because planning and execution are deeply interleaved, we aim to support human-agent collaboration throughout fluid movements between the two phases.
@@
We introduce Cocoa, an interactive system that implements our approach by supporting hybrid human-agent workflows for scientific researchers in a document editor—a common site for research work. This interleaving happens via the following three ways. First, Cocoa supports human-agent co-planning: the agent, when invoked in the document, will first propose a plan of action that is fully interactive and integrated within the document. The user can then edit steps, add new ones, and assign each step to either themselves or the agent. Second, Cocoa supports human-agent co-execution: drawing inspiration from the design of computational notebooks, Cocoa allows the user and the agent to collaboratively complete plan steps, re-executing steps as desired. The user can interactively refine the agent’s intermediate outputs and also manually take over steps themselves to steer the workflow in a more desirable direction. Third, and most importantly, Cocoa interleaves co-planning and co-execution: users can smoothly transition between the two and modify their plans based on outputs from execution, and vice versa.
We then evaluate Cocoa through a lab study with 16 researchers, where we found that interleaving co-planning and co-execution in Cocoa afforded greater agent steerability while maintaining ease of use compared to a strong chat baseline. We finally conducted a 7-day field deployment of Cocoa with 7 researchers, where they found the explicit delegation of work between themselves and the agent, in addition to interleaved co-planning and co-execution, to be valuable in their day-to-day research.
@@
    A formative study with 9 researchers that motivates the design of interactive systems that can flexibly delegate work between the user and agent across task planning and execution.
@@
    Cocoa, an interactive system that interleaves human-agent co-planning and co-execution in a document editor for scientific researchers.
@@
    A lab evaluation of Cocoa with 16 researchers, where we found that participants frequently made use of interleaving co-planning and co-execution when completing research tasks, suggesting that the two are mutually beneficial while enabling more granular steering of the agent.
@@
    A 7-day deployment study of Cocoa with 7 researchers, where participants made use of self- and AI-assigned steps to flexibly delegate human-AI agency and embed their expertise into the workflow.
@@ 2.1. Planning and Interactivity in LLM Agents
Prior work has identified core components in LLM agent architectures to consist of memory, reasoning, planning, and tool use (Yao et al., 2022b; Zhou et al., 2023; Liu et al., 2024; Sumers et al., 2023; Wang et al., 2023; Kim et al., 2024). Central to LLM agents’ longitudinal operations is long-term planning, which demands capable reasoning abilities (Hao et al., 2023) and thus is often implemented by chain-of-thought (CoT) [^1] We use CoT as an umbrella term encompassing chain-of-thought, trees-of-thought (Yao et al., 2023; Zhuang et al., 2023), and related methods. (Wei et al., 2022; Qin et al., 2023; Yao et al., 2022b; OpenAI, 2024b). CoT provides a series of intermediate reasoning steps as exemplars in prompting to boost performance on complex reasoning tasks (Wei et al., 2022; OpenAI, 2024b). CoT is crucial for facilitating task decomposition and therefore LLM agents’ planning capabilities (Qin et al., 2023; Chen et al., 2023). However, even with CoT, critical investigations of LLMs’ abilities to autonomously generate executable plans reveal limited success across diverse domains (Valmeekam et al., 2023). Indeed, although CoT can help improve model planning for tasks with well-defined, objective solutions that provide clear signals for reinforcement learning—e.g., numerical, tabular, and knowledge-based reasoning (Kim et al., 2024)—it may not be so effective in domains that require expert tacit knowledge to navigate ambiguous problem spaces with no single right answer (Hao et al., 2023). Unlike model properties that show empirical improvement through scaling laws, limitations of planning in these domains may not resolve with scale alone as 1) tacit knowledge is not well-documented in training data and is thus difficult for a model to robustly learn (Cheong et al., 2024; Feng et al., 2023; Zhang et al., 2024), and 2) there may be no “correct” workflow for CoT to follow and verify the correctness post-hoc (Jirotka et al., 2013). Scientific research is one such domain (Zhang et al., 2024).
@@
In light of this, recent work has recognized the need to interactively incorporate user feedback for improving LLM agents’ planning capabilities and beyond (Kazemitabaar et al., 2024; Lawley and MacLellan, 2024; Shao et al., 2024; OpenAI, 2025a; Huq et al., 2025; Mozannar et al., 2025; Peng et al., 2025), particularly within scientific discovery (Majumder et al., 2024; Gottweis et al., 2025; Jansen et al., 2025). For example, Lawley and MacLellan (Lawley and MacLellan, 2024) architected an approach where user interaction is used to guide the model in planning for unseen tasks on-the-fly using a hierarchical network of smaller actions. OpenAI’s Operator (OpenAI, 2025a) allows users to “take control” of the agent to perform tasks manually. Yet, interactive techniques in this area are still nascent, despite the new challenges identified for human-agent communication (Bansal et al., 2024). Interfaces for LLM assistants (e.g., AgentGPT (Reworkd, 2024), Devin (Wu, 2024)) have been primarily limited to chat interfaces targeted at monitoring agent activity rather than empowering the user to proactively collaborate with the agent.
An emerging body of work has started to experiment with more interactive techniques for agents. He et al. (He et al., 2025) separated agent planning from execution and found that user involvement during execution can significantly boost agent performance, but poses a high cognitive load. Kazemitabaar et al. (Kazemitabaar et al., 2024) developed interfaces for data analysts to edit an execution plan of an LLM to provide more control points for steering behavior, while Google’s Deep Research (Google, 2025) can re-generate a plan if the user provides feedback via prompting. However, these interactions are corrective rather than collaborative—that is, users would typically engage in these interactions to correct agent behavior post-hoc rather than proactively iterate back-and-forth with the agent at multiple points in the workflow. We do this by introducing affordances for collaborative human-AI planning and execution via a shared representation of tasks and delegation of agency. While our work shares some conceptual similarities with Kazemitabaar et al. (Kazemitabaar et al., 2024), we focus on enabling human-AI co-agency across steps in multi-step plans rather than revealing hidden model reasoning at each step.
@@ 2.2. Computational Notebooks
Donald Knuth’s vision of literate programming interleaves computer programs with natural language documentation (Knuth, 1984). This vision later gave rise to the computational notebook paradigm, which organizes program imperatives, input data, and rich outputs (e.g., tables, visualizations, interactive widgets) into linearly arranged cells (Chattopadhyay et al., 2020; Lau et al., 2020). Computational notebooks are a well-studied paradigm in HCI (e.g., (Choi et al., 2023; Ayobi et al., 2023; Zheng et al., 2022; Lin et al., 2023; Chattopadhyay et al., 2020; O’Leary et al., 2023; Li et al., 2023)), with most prior work focusing on how data scientists leverage them in their workflows.
In our work, we take inspiration from design decisions made for computational notebooks because of the many parallels between workflows in notebooks and mixed-initiative task completion systems (Horvitz, 1999; Grunde-McLaughlin et al., 2023; Wu et al., 2023). For instance, data scientists often break down a high-level data analysis task into executable cells in computational notebooks (Kery et al., 2018; Rule et al., 2018). As they execute each cell and inspect outputs, they can verify the quality and sensibility of outputs before moving forward in their analysis, or go back to modify existing code cells before execution to interleave between programming, execution, and output examination (Kery et al., 2020). Finally, the notebook can serve as documentation and communication of their activities (Li et al., 2023).
By drawing these parallels, we can reveal new opportunities to design interactive workflows and interfaces for LLM agents. Just like how a data scientist may easily add new cells in a notebook or reconfigure existing cells to adapt their analysis plan, a user may interactively edit an agent’s plan of execution to better steer the agent in productive directions. Furthermore, an agentic system may simultaneously be more usable and resource-efficient if a user could direct an agent to iterate on a subtask to improve its output before continuing onto subtasks dependent on that output. This also demands new ways of viewing and editing agent outputs on subtasks, which increases opportunities for human input over existing approaches such as simply logging agent actions (AutoGPT, 2024; Reworkd, 2024; OpenAI Platform, 2024). Our work exploits these parallels to contribute new design patterns for steering agentic systems. We also apply novel extensions of these parallels for human-agent collaboration: agent-proposed plans (i.e, dynamic AI-generated notebook cells) and agent vs. user steps (i.e, AI executes some cells while the user executes others).